OpenStudy (anonymous):

Which is the graph of the line with equation: y + 4 = 2(x – 1)

OpenStudy (anonymous):

y + 4 = 2x -2 y = 2x -6 slope = 2 and y intercept = -6

OpenStudy (anonymous):

y = mx + b m= slope b = y-intercept

OpenStudy (anonymous):

Thanks but, nevermind. :) It timed out. :)

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

Not really funny....... I've been working on it all day..... :(

OpenStudy (anonymous):

oh..sorry

OpenStudy (anonymous):

but you just hav to put the equation into y= mx + b form and graph it

OpenStudy (anonymous):

k...